### v2

# Strong and weak $CP$ tests in sequential decays of polarized $\Sigma^{0}$ hyperons

Strong and weak $CP$ tests in sequential decays of polarized $\Sigma^{0}$ hyperons
M. Ablikim1, M. N. Achasov4,c, P. Adlarson76, O. Afedulidis3, X. C. Ai81, R. Aliberti35, A. Amoroso75A,75C, Q. An72,58,a, Y. Bai57, O. Bakina36, I. Balossino29A, Y. Ban46,h, H.-R. Bao64, V. Batozskaya1,44, K. Begzsuren32, N. Berger35, M. Berlowski44, M. Bertani28A, D. Bettoni29A, F. Bianchi75A,75C, E. Bianco75A,75C, A. Bortone75A,75C, I. Boyko36, R. A. Briere5, A. Brueggemann69, H. Cai77, X. Cai1,58, A. Calcaterra28A, G. F. Cao1,64, N. Cao1,64, S. A. Cetin62A, J. F. Chang1,58, G. R. Che43, G. Chelkov36,b, C. Chen43, C. H. Chen9, Chao Chen55, G. Chen1, H. S. Chen1,64, H. Y. Chen20, M. L. Chen1,58,64, S. J. Chen42, S. L. Chen45, S. M. Chen61, T. Chen1,64, X. R. Chen31,64, X. T. Chen1,64, Y. B. Chen1,58, Y. Q. Chen34, Z. J. Chen25,i, Z. Y. Chen1,64, S. K. Choi10, G. Cibinetto29A, F. Cossio75C, J. J. Cui50, H. L. Dai1,58, J. P. Dai79, A. Dbeyssi18, R.  E. de Boer3, D. Dedovich36, C. Q. Deng73, Z. Y. Deng1, A. Denig35, I. Denysenko36, M. Destefanis75A,75C, F. De Mori75A,75C, B. Ding67,1, X. X. Ding46,h, Y. Ding34, Y. Ding40, J. Dong1,58, L. Y. Dong1,64, M. Y. Dong1,58,64, X. Dong77, M. C. Du1, S. X. Du81, Y. Y. Duan55, Z. H. Duan42, P. Egorov36,b, Y. H. Fan45, J. Fang59, J. Fang1,58, S. S. Fang1,64, W. X. Fang1, Y. Fang1, Y. Q. Fang1,58, R. Farinelli29A, L. Fava75B,75C, F. Feldbauer3, G. Felici28A, C. Q. Feng72,58, J. H. Feng59, Y. T. Feng72,58, M. Fritsch3, C. D. Fu1, J. L. Fu64, Y. W. Fu1,64, H. Gao64, X. B. Gao41, Y. N. Gao46,h, Yang Gao72,58, S. Garbolino75C, I. Garzia29A,29B, L. Ge81, P. T. Ge19, Z. W. Ge42, C. Geng59, E. M. Gersabeck68, A. Gilman70, K. Goetzen13, L. Gong40, W. X. Gong1,58, W. Gradl35, S. Gramigna29A,29B, M. Greco75A,75C, M. H. Gu1,58, Y. T. Gu15, C. Y. Guan1,64, A. Q. Guo31,64, L. B. Guo41, M. J. Guo50, R. P. Guo49, Y. P. Guo12,g, A. Guskov36,b, J. Gutierrez27, K. L. Han64, T. T. Han1, F. Hanisch3, X. Q. Hao19, F. A. Harris66, K. K. He55, K. L. He1,64, F. H. Heinsius3, C. H. Heinz35, Y. K. Heng1,58,64, C. Herold60, T. Holtmann3, P. C. Hong34, G. Y. Hou1,64, X. T. Hou1,64, Y. R. Hou64, Z. L. Hou1, B. Y. Hu59, H. M. Hu1,64, J. F. Hu56,j, S. L. Hu12,g, T. Hu1,58,64, Y. Hu1, Z. M. Hu59, G. S. Huang72,58, K. X. Huang59, L. Q. Huang31,64, X. T. Huang50, Y. P. Huang1, Y. S. Huang59, T. Hussain74, F. Hölzken3, N. Hüsken35, N. in der Wiesche69, J. Jackson27, S. Janchiv32, J. H. Jeong10, Q. Ji1, Q. P. Ji19, W. Ji1,64, X. B. Ji1,64, X. L. Ji1,58, Y. Y. Ji50, X. Q. Jia50, Z. K. Jia72,58, D. Jiang1,64, H. B. Jiang77, P. C. Jiang46,h, S. S. Jiang39, T. J. Jiang16, X. S. Jiang1,58,64, Y. Jiang64, J. B. Jiao50, J. K. Jiao34, Z. Jiao23, S. Jin42, Y. Jin67, M. Q. Jing1,64, X. M. Jing64, T. Johansson76, S. Kabana33, N. Kalantar-Nayestanaki65, X. L. Kang9, X. S. Kang40, M. Kavatsyuk65, B. C. Ke81, V. Khachatryan27, A. Khoukaz69, R. Kiuchi1, O. B. Kolcu62A, B. Kopf3, M. Kuessner3, X. Kui1,64, N.  Kumar26, A. Kupsc44,76, W. Kühn37, J. J. Lane68, L. Lavezzi75A,75C, T. T. Lei72,58, Z. H. Lei72,58, M. Lellmann35, T. Lenz35, C. Li47, C. Li43, C. H. Li39, Cheng Li72,58, D. M. Li81, F. Li1,58, G. Li1, H. B. Li1,64, H. J. Li19, H. N. Li56,j, Hui Li43, J. R. Li61, J. S. Li59, K. Li1, K. L. Li19, L. J. Li1,64, L. K. Li1, Lei Li48, M. H. Li43, P. R. Li38,k,l, Q. M. Li1,64, Q. X. Li50, R. Li17,31, S. X. Li12, T.  Li50, W. D. Li1,64, W. G. Li1,a, X. Li1,64, X. H. Li72,58, X. L. Li50, X. Y. Li1,64, X. Z. Li59, Y. G. Li46,h, Z. J. Li59, Z. Y. Li79, C. Liang42, H. Liang1,64, H. Liang72,58, Y. F. Liang54, Y. T. Liang31,64, G. R. Liao14, Y. P. Liao1,64, J. Libby26, A.  Limphirat60, C. C. Lin55, D. X. Lin31,64, T. Lin1, B. J. Liu1, B. X. Liu77, C. Liu34, C. X. Liu1, F. Liu1, F. H. Liu53, Feng Liu6, G. M. Liu56,j, H. Liu38,k,l, H. B. Liu15, H. H. Liu1, H. M. Liu1,64, Huihui Liu21, J. B. Liu72,58, J. Y. Liu1,64, K. Liu38,k,l, K. Y. Liu40, Ke Liu22, L. Liu72,58, L. C. Liu43, Lu Liu43, M. H. Liu12,g, P. L. Liu1, Q. Liu64, S. B. Liu72,58, T. Liu12,g, W. K. Liu43, W. M. Liu72,58, X. Liu39, X. Liu38,k,l, Y. Liu81, Y. Liu38,k,l, Y. B. Liu43, Z. A. Liu1,58,64, Z. D. Liu9, Z. Q. Liu50, X. C. Lou1,58,64, F. X. Lu59, H. J. Lu23, J. G. Lu1,58, X. L. Lu1, Y. Lu7, Y. P. Lu1,58, Z. H. Lu1,64, C. L. Luo41, J. R. Luo59, M. X. Luo80, T. Luo12,g, X. L. Luo1,58, X. R. Lyu64, Y. F. Lyu43, F. C. Ma40, H. Ma79, H. L. Ma1, J. L. Ma1,64, L. L. Ma50, L. R. Ma67, M. M. Ma1,64, Q. M. Ma1, R. Q. Ma1,64, T. Ma72,58, X. T. Ma1,64, X. Y. Ma1,58, Y. Ma46,h, Y. M. Ma31, F. E. Maas18, M. Maggiora75A,75C, S. Malde70, Q. A. Malik74, Y. J. Mao46,h, Z. P. Mao1, S. Marcello75A,75C, Z. X. Meng67, J. G. Messchendorp13,65, G. Mezzadri29A, H. Miao1,64, T. J. Min42, R. E. Mitchell27, X. H. Mo1,58,64, B. Moses27, N. Yu. Muchnoi4,c, J. Muskalla35, Y. Nefedov36, F. Nerling18,e, L. S. Nie20, I. B. Nikolaev4,c, Z. Ning1,58, S. Nisar11,m, Q. L. Niu38,k,l, W. D. Niu55, Y. Niu 50, S. L. Olsen64, Q. Ouyang1,58,64, S. Pacetti28B,28C, X. Pan55, Y. Pan57, A.  Pathak34, Y. P. Pei72,58, M. Pelizaeus3, H. P. Peng72,58, Y. Y. Peng38,k,l, K. Peters13,e, J. L. Ping41, R. G. Ping1,64, S. Plura35, V. Prasad33, F. Z. Qi1, H. Qi72,58, H. R. Qi61, M. Qi42, T. Y. Qi12,g, S. Qian1,58, W. B. Qian64, C. F. Qiao64, X. K. Qiao81, J. J. Qin73, L. Q. Qin14, L. Y. Qin72,58, X. P. Qin12,g, X. S. Qin50, Z. H. Qin1,58, J. F. Qiu1, Z. H. Qu73, C. F. Redmer35, K. J. Ren39, A. Rivetti75C, M. Rolo75C, G. Rong1,64, Ch. Rosner18, S. N. Ruan43, N. Salone44, A. Sarantsev36,d, Y. Schelhaas35, K. Schoenning76, M. Scodeggio29A, K. Y. Shan12,g, W. Shan24, X. Y. Shan72,58, Z. J. Shang38,k,l, J. F. Shangguan16, L. G. Shao1,64, M. Shao72,58, C. P. Shen12,g, H. F. Shen1,8, W. H. Shen64, X. Y. Shen1,64, B. A. Shi64, H. Shi72,58, H. C. Shi72,58, J. L. Shi12,g, J. Y. Shi1, Q. Q. Shi55, S. Y. Shi73, X. Shi1,58, J. J. Song19, T. Z. Song59, W. M. Song34,1, Y.  J. Song12,g, Y. X. Song46,h,n, S. Sosio75A,75C, S. Spataro75A,75C, F. Stieler35, S. S Su40, Y. J. Su64, G. B. Sun77, G. X. Sun1, H. Sun64, H. K. Sun1, J. F. Sun19, K. Sun61, L. Sun77, S. S. Sun1,64, T. Sun51,f, W. Y. Sun34, Y. Sun9, Y. J. Sun72,58, Y. Z. Sun1, Z. Q. Sun1,64, Z. T. Sun50, C. J. Tang54, G. Y. Tang1, J. Tang59, M. Tang72,58, Y. A. Tang77, L. Y. Tao73, Q. T. Tao25,i, M. Tat70, J. X. Teng72,58, V. Thoren76, W. H. Tian59, Y. Tian31,64, Z. F. Tian77, I. Uman62B, Y. Wan55, S. J. Wang 50, B. Wang1, B. L. Wang64, Bo Wang72,58, D. Y. Wang46,h, F. Wang73, H. J. Wang38,k,l, J. J. Wang77, J. P. Wang 50, K. Wang1,58, L. L. Wang1, M. Wang50, N. Y. Wang64, S. Wang12,g, S. Wang38,k,l, T.  Wang12,g, T. J. Wang43, W.  Wang73, W. Wang59, W. P. Wang35,58,72,o, X. Wang46,h, X. F. Wang38,k,l, X. J. Wang39, X. L. Wang12,g, X. N. Wang1, Y. Wang61, Y. D. Wang45, Y. F. Wang1,58,64, Y. L. Wang19, Y. N. Wang45, Y. Q. Wang1, Yaqian Wang17, Yi Wang61, Z. Wang1,58, Z. L.  Wang73, Z. Y. Wang1,64, Ziyi Wang64, D. H. Wei14, F. Weidner69, S. P. Wen1, Y. R. Wen39, U. Wiedner3, G. Wilkinson70, M. Wolke76, L. Wollenberg3, C. Wu39, J. F. Wu1,8, L. H. Wu1, L. J. Wu1,64, X. Wu12,g, X. H. Wu34, Y. Wu72,58, Y. H. Wu55, Y. J. Wu31, Z. Wu1,58, L. Xia72,58, X. M. Xian39, B. H. Xiang1,64, T. Xiang46,h, D. Xiao38,k,l, G. Y. Xiao42, S. Y. Xiao1, Y.  L. Xiao12,g, Z. J. Xiao41, C. Xie42, X. H. Xie46,h, Y. Xie50, Y. G. Xie1,58, Y. H. Xie6, Z. P. Xie72,58, T. Y. Xing1,64, C. F. Xu1,64, C. J. Xu59, G. F. Xu1, H. Y. Xu67,2,p, M. Xu72,58, Q. J. Xu16, Q. N. Xu30, W. Xu1, W. L. Xu67, X. P. Xu55, Y. Xu40, Y. C. Xu78, Z. S. Xu64, F. Yan12,g, L. Yan12,g, W. B. Yan72,58, W. C. Yan81, X. Q. Yan1,64, H. J. Yang51,f, H. L. Yang34, H. X. Yang1, T. Yang1, Y. Yang12,g, Y. F. Yang1,64, Y. F. Yang43, Y. X. Yang1,64, Z. W. Yang38,k,l, Z. P. Yao50, M. Ye1,58, M. H. Ye8, J. H. Yin1, Junhao Yin43, Z. Y. You59, B. X. Yu1,58,64, C. X. Yu43, G. Yu1,64, J. S. Yu25,i, M. C. Yu40, T. Yu73, X. D. Yu46,h, Y. C. Yu81, C. Z. Yuan1,64, J. Yuan34, J. Yuan45, L. Yuan2, S. C. Yuan1,64, Y. Yuan1,64, Z. Y. Yuan59, C. X. Yue39, A. A. Zafar74, F. R. Zeng50, S. H. Zeng63A,63B,63C,63D, X. Zeng12,g, Y. Zeng25,i, Y. J. Zeng59, Y. J. Zeng1,64, X. Y. Zhai34, Y. C. Zhai50, Y. H. Zhan59, A. Q. Zhang1,64, B. L. Zhang1,64, B. X. Zhang1, D. H. Zhang43, G. Y. Zhang19, H. Zhang81, H. Zhang72,58, H. C. Zhang1,58,64, H. H. Zhang59, H. H. Zhang34, H. Q. Zhang1,58,64, H. R. Zhang72,58, H. Y. Zhang1,58, J. Zhang81, J. Zhang59, J. J. Zhang52, J. L. Zhang20, J. Q. Zhang41, J. S. Zhang12,g, J. W. Zhang1,58,64, J. X. Zhang38,k,l, J. Y. Zhang1, J. Z. Zhang1,64, Jianyu Zhang64, L. M. Zhang61, Lei Zhang42, P. Zhang1,64, Q. Y. Zhang34, R. Y. Zhang38,k,l, S. H. Zhang1,64, Shulei Zhang25,i, X. D. Zhang45, X. M. Zhang1, X. Y Zhang40, X. Y. Zhang50, Y.  Zhang73, Y. Zhang1, Y.  T. Zhang81, Y. H. Zhang1,58, Y. M. Zhang39, Yan Zhang72,58, Z. D. Zhang1, Z. H. Zhang1, Z. L. Zhang34, Z. Y. Zhang77, Z. Y. Zhang43, Z. Z.  Zhang45, G. Zhao1, J. Y. Zhao1,64, J. Z. Zhao1,58, L. Zhao1, Lei Zhao72,58, M. G. Zhao43, N. Zhao79, R. P. Zhao64, S. J. Zhao81, Y. B. Zhao1,58, Y. X. Zhao31,64, Z. G. Zhao72,58, A. Zhemchugov36,b, B. Zheng73, B. M. Zheng34, J. P. Zheng1,58, W. J. Zheng1,64, Y. H. Zheng64, B. Zhong41, X. Zhong59, H.  Zhou50, J. Y. Zhou34, L. P. Zhou1,64, S.  Zhou6, X. Zhou77, X. K. Zhou6, X. R. Zhou72,58, X. Y. Zhou39, Y. Z. Zhou12,g, Z. C. Zhou20, A. N. Zhu64, J. Zhu43, K. Zhu1, K. J. Zhu1,58,64, K. S. Zhu12,g, L. Zhu34, L. X. Zhu64, S. H. Zhu71, T. J. Zhu12,g, W. D. Zhu41, Y. C. Zhu72,58, Z. A. Zhu1,64, J. H. Zou1, J. Zu72,58
(BESIII Collaboration)
1 Institute of High Energy Physics, Beijing 100049, People’s Republic of China
2 Beihang University, Beijing 100191, People’s Republic of China
3 Bochum Ruhr-University, D-44780 Bochum, Germany
4 Budker Institute of Nuclear Physics SB RAS (BINP), Novosibirsk 630090, Russia
5 Carnegie Mellon University, Pittsburgh, Pennsylvania 15213, USA
6 Central China Normal University, Wuhan 430079, People’s Republic of China
7 Central South University, Changsha 410083, People’s Republic of China
8 China Center of Advanced Science and Technology, Beijing 100190, People’s Republic of China
9 China University of Geosciences, Wuhan 430074, People’s Republic of China
10 Chung-Ang University, Seoul, 06974, Republic of Korea
11 COMSATS University Islamabad, Lahore Campus, Defence Road, Off Raiwind Road, 54000 Lahore, Pakistan
12 Fudan University, Shanghai 200433, People’s Republic of China
13 GSI Helmholtzcentre for Heavy Ion Research GmbH, D-64291 Darmstadt, Germany
14 Guangxi Normal University, Guilin 541004, People’s Republic of China
15 Guangxi University, Nanning 530004, People’s Republic of China
16 Hangzhou Normal University, Hangzhou 310036, People’s Republic of China
17 Hebei University, Baoding 071002, People’s Republic of China
18 Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
19 Henan Normal University, Xinxiang 453007, People’s Republic of China
20 Henan University, Kaifeng 475004, People’s Republic of China
21 Henan University of Science and Technology, Luoyang 471003, People’s Republic of China
22 Henan University of Technology, Zhengzhou 450001, People’s Republic of China
23 Huangshan College, Huangshan 245000, People’s Republic of China
24 Hunan Normal University, Changsha 410081, People’s Republic of China
25 Hunan University, Changsha 410082, People’s Republic of China
26 Indian Institute of Technology Madras, Chennai 600036, India
27 Indiana University, Bloomington, Indiana 47405, USA
28 INFN Laboratori Nazionali di Frascati , (A)INFN Laboratori Nazionali di Frascati, I-00044, Frascati, Italy; (B)INFN Sezione di Perugia, I-06100, Perugia, Italy; (C)University of Perugia, I-06100, Perugia, Italy
29 INFN Sezione di Ferrara, (A)INFN Sezione di Ferrara, I-44122, Ferrara, Italy; (B)University of Ferrara, I-44122, Ferrara, Italy
30 Inner Mongolia University, Hohhot 010021, People’s Republic of China
31 Institute of Modern Physics, Lanzhou 730000, People’s Republic of China
32 Institute of Physics and Technology, Peace Avenue 54B, Ulaanbaatar 13330, Mongolia
33 Instituto de Alta Investigación, Universidad de Tarapacá, Casilla 7D, Arica 1000000, Chile
34 Jilin University, Changchun 130012, People’s Republic of China
35 Johannes Gutenberg University of Mainz, Johann-Joachim-Becher-Weg 45, D-55099 Mainz, Germany
36 Joint Institute for Nuclear Research, 141980 Dubna, Moscow region, Russia
37 Justus-Liebig-Universitaet Giessen, II. Physikalisches Institut, Heinrich-Buff-Ring 16, D-35392 Giessen, Germany
38 Lanzhou University, Lanzhou 730000, People’s Republic of China
39 Liaoning Normal University, Dalian 116029, People’s Republic of China
40 Liaoning University, Shenyang 110036, People’s Republic of China
41 Nanjing Normal University, Nanjing 210023, People’s Republic of China
42 Nanjing University, Nanjing 210093, People’s Republic of China
43 Nankai University, Tianjin 300071, People’s Republic of China
44 National Centre for Nuclear Research, Warsaw 02-093, Poland
45 North China Electric Power University, Beijing 102206, People’s Republic of China
46 Peking University, Beijing 100871, People’s Republic of China
47 Qufu Normal University, Qufu 273165, People’s Republic of China
48 Renmin University of China, Beijing 100872, People’s Republic of China
49 Shandong Normal University, Jinan 250014, People’s Republic of China
50 Shandong University, Jinan 250100, People’s Republic of China
51 Shanghai Jiao Tong University, Shanghai 200240, People’s Republic of China
52 Shanxi Normal University, Linfen 041004, People’s Republic of China
53 Shanxi University, Taiyuan 030006, People’s Republic of China
54 Sichuan University, Chengdu 610064, People’s Republic of China
55 Soochow University, Suzhou 215006, People’s Republic of China
56 South China Normal University, Guangzhou 510006, People’s Republic of China
57 Southeast University, Nanjing 211100, People’s Republic of China
58 State Key Laboratory of Particle Detection and Electronics, Beijing 100049, Hefei 230026, People’s Republic of China
59 Sun Yat-Sen University, Guangzhou 510275, People’s Republic of China
60 Suranaree University of Technology, University Avenue 111, Nakhon Ratchasima 30000, Thailand
61 Tsinghua University, Beijing 100084, People’s Republic of China
62 Turkish Accelerator Center Particle Factory Group, (A)Istinye University, 34010, Istanbul, Turkey; (B)Near East University, Nicosia, North Cyprus, 99138, Mersin 10, Turkey
63 University of Bristol, (A)H H Wills Physics Laboratory; (B)Tyndall Avenue; (C)Bristol; (D)BS8 1TL
64 University of Chinese Academy of Sciences, Beijing 100049, People’s Republic of China
65 University of Groningen, NL-9747 AA Groningen, The Netherlands
66 University of Hawaii, Honolulu, Hawaii 96822, USA
67 University of Jinan, Jinan 250022, People’s Republic of China
68 University of Manchester, Oxford Road, Manchester, M13 9PL, United Kingdom
69 University of Muenster, Wilhelm-Klemm-Strasse 9, 48149 Muenster, Germany
70 University of Oxford, Keble Road, Oxford OX13RH, United Kingdom
71 University of Science and Technology Liaoning, Anshan 114051, People’s Republic of China
72 University of Science and Technology of China, Hefei 230026, People’s Republic of China
73 University of South China, Hengyang 421001, People’s Republic of China
74 University of the Punjab, Lahore-54590, Pakistan
75 University of Turin and INFN, (A)University of Turin, I-10125, Turin, Italy; (B)University of Eastern Piedmont, I-15121, Alessandria, Italy; (C)INFN, I-10125, Turin, Italy
76 Uppsala University, Box 516, SE-75120 Uppsala, Sweden
77 Wuhan University, Wuhan 430072, People’s Republic of China
78 Yantai University, Yantai 264005, People’s Republic of China
79 Yunnan University, Kunming 650500, People’s Republic of China
80 Zhejiang University, Hangzhou 310027, People’s Republic of China
81 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
a Deceased
b Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
c Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
d Also at the NRC ”Kurchatov Institute”, PNPI, 188300, Gatchina, Russia
e Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
f Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
g Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
h Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
i Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
j Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
k Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
m Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
n Also at Ecole Polytechnique Federale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
o Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
p Also at School of Physics, Beihang University, Beijing 100191 , China
Abstract
The $J/\psi,\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ processes and subsequent decays are studied using the world’s largest $J/\psi$ and $\psi(3686)$ data samples collected with the BESIII detector. The parity-violating decay parameters of the decays $\Sigma^{0}\to\Lambda\gamma$ and $\bar{\Sigma}^{0}\to\bar{\Lambda}\gamma$, $\alpha_{\Sigma^{0}}=-0.0017\pm 0.0021\pm 0.0018$ and $\bar{\alpha}_{\Sigma^{0}}=0.0021\pm 0.0020\pm 0.0022$, are measured for the first time. The strong-$CP$ symmetry is tested in the decays of the $\Sigma^{0}$ hyperons for the first time by measuring the asymmetry $A^{\Sigma}_{CP}=\alpha_{\Sigma^{0}}+\bar{\alpha}_{\Sigma^{0}}=(0.4\pm 2.9\pm 1.3)\times 10^{-3}$. The weak-$CP$ test is performed in the subsequent decays of their daughter particles $\Lambda$ and $\bar{\Lambda}$. Also for the first time, the transverse polarizations of the $\Sigma^{0}$ hyperons in $J/\psi$ and $\psi(3686)$ decays are observed with opposite directions, and the ratios between the S-wave and D-wave contributions of the $J/\psi,\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ decays are obtained. These results are crucial to understand the decay dynamics of the charmonium states and the production mechanism of the $\Sigma^{0}-\bar{\Sigma}^{0}$ pairs.
The question of why our universe consists of matter, but almost no antimatter, has puzzled the scientific community for more than half a century and is to this day subjected to intensive research Ablikim et al. (2022a, b, 2023a); Aaij et al. (2024); Guan et al. ; Aaij et al. . One of the long-standing explanations is baryogenesis Sakharov (1967), i.e. that the matter abundance is generated dynamically. Among the necessary criteria for this to be possible is the existence of processes that violate charge-conjugation and parity ($CP$) conservation. The Standard Model (SM) of particle physics allows for tiny signals of $CP$ violation ($CPV$) in weak interactions, in line with experimental observations in meson decays Christenson et al. (1964); Aubert et al. (2001); Abe et al. (2001); Aaij et al. (2019). The SM should in principle also allow for $CP$ violating strong processes, which would be manifest in e.g. a non-zero neutron electric dipole moment (EDM) Ramsey (1982); Fortson et al. (2003); Luders (1957). However, the extremely tiny experimental upper limit of the neutron EDM Abel et al. (2020) suggests that $CPV$ in strong interactions is unnaturally small. The SM itself does not provide an answer to the smallness of strong $CPV$, which calls for further investigations. The radiative decay $\Sigma^{0}\to\Lambda\gamma$ could in principle offer an opportunity to study the interference between a parity-conserving (from the magnetic transition moment) and a parity-violating amplitude (from the electric dipole transition moment, EDTM). The EDTM of this decay is related to the neutron EDM via SU(3) flavor symmetry Gell-Mann (1961); Nair et al. (2019). This can be explored by measuring the parity-violating decay parameters $\alpha_{\Sigma^{0}}$, and $\bar{\alpha}_{\Sigma^{0}}$ for the charge conjugate decay. In the SM, $\alpha_{\Sigma^{0}}$ and $\bar{\alpha}_{\Sigma^{0}}$ are predicted to be very small Nair et al. (2019), nonzero measurements of these parameters at current or near-future experimental sensitivity levels would be indications of parity violation beyond the standard model. Based on the values of these parameters, the strong-$CP$ symmetry can be tested by measuring the asymmetry $A^{\Sigma}_{CP}=\alpha_{\Sigma^{0}}+\bar{\alpha}_{\Sigma^{0}}$, finding a non-vanishing $A^{\Sigma}_{CP}$ would point to signals of $CPV$ beyond the SM Nair et al. (2019).
The large yield of quantum entangled $\Sigma^{0}\bar{\Sigma}^{0}$ pairs at BESIII Li (2017) enables a pioneering test of strong-$CP$ symmetry in the $\Sigma^{0}$ hyperon decays. In addition, the subsequent decays of the daughter particles $\Lambda$ and $\bar{\Lambda}$ provide an independent measurement of the Λ/$\Lambda/\bar{\Lambda}$ decay parameter $\alpha_{\Lambda}/\bar{\alpha}_{\Lambda}$, which have undergone much scrutiny in recent years due to the large discrepancy observed between old and new data Ablikim et al. (2019, 2022b, 2022a, 2023a); Astbury et al. (1975); Cleland et al. (1972); Overseth and Roth (1967); Dauber et al. (1969), and allow a test of weak-$CP$ symmetry. Finally, the data collected with different vector charmonia will help elucidate their decay dynamics.
The decay of a charmonium ($\psi$) into a hyperon-antihyperon pair can be completely described by two parameters, $\alpha_{\psi}$ and $\Delta\Phi_{\psi}$ Fäldt and Kupsc (2017). The former describes the angular distribution of the hyperon-antihyperon pair, while the latter is related to the hyperon polarization,
$P_{y}(\cos\theta_{Y})=\frac{\sqrt{1-\alpha^{2}_{\psi}}\sin\Delta\Phi_{\psi}\cos\theta_{Y}\sin\theta_{Y}}{1+\alpha_{\psi}\cos^{2}\theta_{Y}},$ (1)
where $\theta_{Y}$ is the angle between the momenta of $e^{+}$ and the hyperon in the $e^{+}e^{-}$ center-of-mass (C.M.) system. The hyperon polarization would manifest itself by a non-zero $\Delta\Phi_{\psi}$.
In recent years, the BESIII collaboration has published a series of studies Ablikim et al. (2019, 2020a, 2021, 2022a, 2022b, 2023a, 2022c, 2023b, 2023c) about hyperon polarization. An intriguing result was reported in Ref. Ablikim et al. (2020a), where the directions of the $\Sigma^{+}/\bar{\Sigma}^{-}$ polarizations were observed to be opposite in $J/\psi$ and $\psi(3686)$ decays; this phenomenon, however, is not observed in $J/\psi\to\Xi^{-}\bar{\Xi}^{+}/\Xi^{0}\bar{\Xi}^{0}$ and $\psi(3686)\to\Xi^{-}\bar{\Xi}^{+}/\Xi^{0}\bar{\Xi}^{0}$ decays Ablikim et al. (2022a, c, 2023b, 2023a, a). Until now, there is no interpretation of these results, therefore, more experimental measurements of the hyperon polarization are highly desirable.
The analysis presented in this Letter is based on samples of $(10087\pm 44)\times 10^{6}$ $J/\psi$ and $(2712\pm 14)\times 10^{6}$ $\psi(3686)$ events Ablikim et al. (2022d, ) collected with the BESIII detector at the BEPCII collider. Details about BEPCII and BESIII can be found in Refs. Ablikim et al. (2010, 2020b); Yu et al. (2016); Huang et al. (2022); Li et al. (2017); Guo et al. (2017); Cao et al. (2020). Simulated data samples produced with geant4-based Agostinelli et al. (2003) Monte Carlo (MC) software, which includes the geometric description of the BESIII detector and the detector response, are used to determine detection efficiencies and estimate backgrounds. The simulations model the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc Jadach et al. (2001). Inclusive MC samples of $J/\psi$ and $\psi(3686)$ resonances are produced, in which the known decay modes are modeled with evtgen Lange (2001); *Ping2008, and the remaining unknown decays are modeled with lundcharm Chen et al. (2000); *Yang2014a. For the signal process, $e^{+}e^{-}\to\psi\to\Sigma^{0}\bar{\Sigma}^{0},\Sigma^{0}\to\Lambda(\to p\pi^{-})\gamma,\bar{\Sigma}^{0}\to\bar{\Lambda}(\to\bar{p}\pi^{+})\gamma$, two different MC samples are used. One is generated, with the uniform distribution of the kinematic momenta of the final states, to obtain the efficiency correction used in the fit to determine the parameters, and the other is generated according to the joint angular distribution for the signal process with the parameters obtained by this analysis (signal MC).
The $\Lambda$ and $\bar{\Lambda}$ hyperons are reconstructed from their dominant decay modes, $\Lambda\to p\pi^{-}$ and $\bar{\Lambda}\to\bar{p}\pi^{+}$. Charged tracks detected in the multilayer drift chamber (MDC) are required to be within a polar angle ($\theta$) range of $|\rm{cos\theta}|<0.93$, where $\theta$ is defined with respect to the $z$-axis, which is the symmetry axis of the MDC. Events with at least four charged tracks are retained. Tracks with momentum larger than $0.45$ GeV/$c$ for the $J/\psi$ dataset and $0.55$ GeV/$c$ for the $\psi(3686)$ dataset are considered as proton candidates, otherwise as pion candidates. There is no further particle identification requirement. Secondary vertex fits Xu et al. (2009) are performed for all combinations with oppositely-charged proton and pion candidates, and the combinations passing the vertex fit successfully and with invariant masses within a range of $[1.111,1.120]$ GeV/$c^{2}$ are regarded as the $\Lambda$ and $\bar{\Lambda}$ candidates.
Photon candidates are identified using showers in the electromagnetic calorimeter (EMC). The deposited energy of each shower must be more than 25 MeV in the barrel region($|\cos\theta|<0.80$) and more than 50 MeV in the end-cap region ($0.920.86<|\cos\theta|<$ 0.92). To suppress electronic noise and showers unrelated to the event, the difference between the EMC time of the photon candidate and the event start time is required to be within [0, 700] ns.
Events with at least one $\Lambda$ candidate, one $\bar{\Lambda}$ candidate, and two photons are considered for further analysis. A four-constraint (4C) kinematic fit is performed under the $\Lambda\bar{\Lambda}\gamma\gamma$ hypothesis by constraining their total reconstructed four-momentum to that of the initial $\psi$. If there is more than one $\Lambda\bar{\Lambda}\gamma\gamma$ combination, the one with the smallest $\chi^{2}$ of the 4C fit ($\chi^{2}_{4\rm C}$) is selected, and $\chi^{2}_{4\rm C}<100$ is required. Since there are two $\gamma$ candidates ($\gamma_{1}$, $\gamma_{2}$) per event, there will be two combinations of Λ($\Lambda(\bar{\Lambda})$ and $\gamma$. Therefore, a six-constraint (6C) kinematic fit is performed under the $\Lambda\bar{\Lambda}\gamma\gamma$ hypothesis by additionally constraining the invariant masses of $\Lambda\gamma_{1}$ $(m_{\Lambda\gamma_{1}})$ and $\bar{\Lambda}\gamma_{2}$ $(m_{\bar{\Lambda}\gamma_{2}})$ to the known masses Workman et al. (2022) of $\Sigma^{0}$ and $\bar{\Sigma}^{0}$. The combination which minimizes $\chi^{2}$ of the 6C fit ($\chi^{2}_{6C}$) is kept, and there is no further requirement on the $\chi^{2}_{6C}$. Finally, the $m_{\Lambda\gamma_{1}}$ and $m_{\bar{\Lambda}\gamma_{2}}$ masses from the 4C fit are required to be within $[1.178,1.206]$ GeV/$c^{2}$. After applying all the event selection criteria, the final event samples in the signal region contain $1083676$ events for the $J/\psi$ decay and $51837$ events for the $\psi(3686)$ decay.
The backgrounds in this analysis are divided into three categories: background channels which have a $\Sigma^{0}\bar{\Sigma}^{0}$ pair (type 1), channels that have one $\Sigma^{0}$ or one $\bar{\Sigma}^{0}$ (type 2), and ones containing no $\Sigma^{0}$ or $\bar{\Sigma}^{0}$ (type 3). Inclusive MC samples of 10 billion $J/\psi$ events and 2.7 billion $\psi(3686)$ events are used for studying background type 1. After applying the same selection criteria as for data, the only peaking background channel is found to be $J/\psi\to\gamma\eta_{c},\eta_{c}\to\Sigma^{0}\bar{\Sigma}^{0}$. An exclusive MC simulation of this process is carried out, and the corresponding number of events from this channel in data is estimated to be $391\pm 145$, which is only about 0.036% of the total selected candidates in the $J/\psi$ data sample and can thus be neglected. The later two types of backgrounds are estimated with the two-dimensional sideband regions of the distribution of $m_{\bar{\Lambda}\gamma_{2}}$ versus $m_{\Lambda\gamma_{1}}$ using the same method as in Ref. Ablikim et al. (2020a). The lower and upper sideband regions are defined as $m_{\Lambda\gamma_{1}/\bar{\Lambda}\gamma_{2}}\in(1.138,1.166)$ GeV/$c^{2}$ and $m_{\Lambda\gamma_{1}/\bar{\Lambda}\gamma_{2}}\in(1.218,1.246)$ GeV/$c^{2}$, respectively. The background levels are found to be 0.7% for $J/\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ and 1.5% for $\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$.
Following the formulation in Refs. Dubnickova et al. (1996); Gakh and Tomasi-Gustafsson (2006); Czyz et al. (2007); Commins and Bucksbaum (1983); Nair et al. (2019); Li and Ma (2019), the joint angular distribution of the full decay chain $e^{+}e^{-}\to\psi\to\Sigma^{0}\bar{\Sigma}^{0},\Sigma^{0}\to\Lambda(\to p\pi^{-})\gamma,\bar{\Sigma}^{0}\to\bar{\Lambda}(\to\bar{p}\pi^{+})\gamma$ is obtained,
$\mathcal{W}(\vec{\zeta},\vec{\omega})\propto(1-\alpha_{\Lambda}\alpha_{\Sigma^{0}}\cos\theta_{p})(1-\bar{\alpha}_{\Lambda}\bar{\alpha}_{\Sigma^{0}}\cos\theta_{\bar{p}})\times$
$\{1+\alpha_{\psi}\cos^{2}\theta_{\Sigma}+\sqrt{1-\alpha^{2}_{\psi}}\sin\theta_{\Sigma}\cos\theta_{\Sigma}\cdot$
$[\beta_{\gamma}\bar{\beta}_{\gamma}F_{1}-(\beta_{\gamma}F_{2}-\bar{\beta}_{\gamma}F_{3})]+$
$\beta_{\gamma}\bar{\beta}_{\gamma}[\alpha_{\psi}F_{4}+F_{5}-(\alpha_{\psi}+\cos^{2}\theta_{\Sigma})\cos\theta_{\Lambda}\cos\theta_{\bar{\Lambda}}]\}.$ (2)
Here, $\vec{\omega}$ represents the six parameters of interest, $\alpha_{\psi}$, $\Delta\Phi_{\psi}$, $\alpha_{\Sigma^{0}}$, $\bar{\alpha}_{\Sigma^{0}}$, $\alpha_{\Lambda}$, $\bar{\alpha}_{\Lambda}$; $\vec{\zeta}$ stands for seven helicity angles, $\theta_{\Sigma}$, $\theta_{\Lambda}$, $\varphi_{\Lambda}$, $\theta_{\bar{\Lambda}}$, $\varphi_{\bar{\Lambda}}$, $\theta_{p}$, and $\theta_{\bar{p}}$. For simplicity, the symbols $\beta_{\gamma}\equiv\frac{\alpha_{\Sigma^{0}}-\alpha_{\Lambda}\cos\theta_{p}}{1-\alpha_{\Lambda}\alpha_{\Sigma^{0}}\cos\theta_{p}}$ and $\bar{\beta}_{\gamma}\equiv\frac{\bar{\alpha}_{\Sigma^{0}}-\bar{\alpha}_{\Lambda}\cos\theta_{\bar{p}}}{1-\bar{\alpha}_{\Lambda}\bar{\alpha}_{\Sigma^{0}}\cos\theta_{\bar{p}}}$ are introduced, and $F_{1}-F_{5}$ are defined below,
$F_{1}=\cos\Delta\Phi_{\psi}(\sin\theta_{\Lambda}\cos\varphi_{\Lambda}\cos\theta_{\bar{\Lambda}}-\sin\theta_{\bar{\Lambda}}\cos\varphi_{\bar{\Lambda}}\cos\theta_{\Lambda}),$
$F_{2}=\sin\Delta\Phi_{\psi}\sin\theta_{\Lambda}\sin\varphi_{\Lambda},$
$F_{3}=\sin\Delta\Phi_{\psi}\sin\theta_{\bar{\Lambda}}\sin\varphi_{\bar{\Lambda}},$
$F_{4}=\sin^{2}\theta_{\Sigma}\sin\theta_{\Lambda}\sin\theta_{\bar{\Lambda}}\sin\varphi_{\Lambda}\sin\varphi_{\bar{\Lambda}},$
$F_{5}=\sin^{2}\theta_{\Sigma}\sin\theta_{\Lambda}\sin\theta_{\bar{\Lambda}}\cos\varphi_{\Lambda}\cos\varphi_{\bar{\Lambda}}.$ (3)
The helicity angles are constructed as illustrated in Fig. 1, and the corresponding angles of the anti-particle decay sequence are obtained analogously. To validate the formula, a check is made by setting the parameters $\alpha_{\Sigma^{0}}$ and $\bar{\alpha}_{\Sigma^{0}}$ in $\mathcal{W}(\vec{\zeta},\vec{\omega})$ to zero, and then the formalism is found to be equivalent to the covariant formalism given in Ref. Fäldt and Schönning (2020).
Figure 1: The definitions of the helicity angles. The polar angle $\theta_{\Sigma}$ is the angle between the $\Sigma^{0}$ momentum and the $e^{+}$ beam direction in the $e^{+}e^{-}$ C.M., where the $\hat{z}$ axis is defined along the $e^{+}$ momentum. $\theta_{\Lambda}$ and $\varphi_{\Lambda}$ are the polar and azimuthal angles of the $\Lambda$ momentum direction in the $\Sigma^{0}$ rest frame, where $\hat{z}_{\Sigma^{0}}$ is defined along the $\Sigma^{0}$ momentum direction in the $e^{+}e^{-}$ C.M., and $\hat{y}_{\Sigma^{0}}$ is defined by $\hat{z}\times\hat{z}_{\Sigma^{0}}$. The angles $\theta_{p}$ and $\varphi_{p}$ are the polar and azimuthal angles of the proton momentum direction in the $\Lambda$ rest frame, where $\hat{z}_{\Lambda}$ is defined along the $\Lambda$ momentum in the $\Sigma^{0}$ rest frame, and $\hat{y}_{\Lambda}$ is along $\hat{z}_{\Sigma^{0}}\times\hat{z}_{\Lambda}$.
An unbinned maximum likelihood fit is performed in the seven angular dimensions $\vec{\zeta}$ by simultaneously fitting both the $J/\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ and $\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ datasets to determine the parameters $\vec{\omega}$. During the fitting process, the background events estimated from the data sidebands are included, and their log-likelihood is subtracted from the data based on the normalized weights. The numerical results are summarized in Table 2, together with the $CP$ asymmetries $A^{\Sigma}_{CP}$ and $A^{\Lambda}_{CP}=(\alpha_{\Lambda}+\bar{\alpha}_{\Lambda})/(\alpha_{\Lambda}-\bar{\alpha}_{\Lambda})$. The correlation coefficient matrix of the fitted parameters is shown in Tab. 1.
Table 1: Full correlation coefficient matrix of the fitted parameters.
Coefficients $\alpha_{J/\psi}$ $\Delta\Phi_{J/\psi}$ $\alpha_{\psi(3686)}$ $\Delta\Phi_{\psi(3686)}$ $\alpha_{\Sigma^{0}}$ $\bar{\alpha}_{\Sigma^{0}}$ $\alpha_{\Lambda}$ $\bar{\alpha}_{\Lambda}$
$\alpha_{J/\psi}$ 1.000 -0.016 -0.012 0.012 -0.002 -0.002 0.018 -0.013
$\Delta\Phi_{J/\psi}$0.016 1.000 0.006 -0.007 0.004 0.000 0.011 0.027
$\alpha_{\psi(3686)}$0.012 0.006 1.000 -0.451 0.007 -0.007 -0.011 0.004
$\Delta\Phi_{\psi(3686)}$ 0.012 -0.007 -0.451 1.000 -0.001 -0.001 -0.051 -0.067
$\alpha_{\Sigma^{0}}$0.002 0.004 0.007 -0.001 1.000 -0.009 -0.083 -0.083
$\bar{\alpha}_{\Sigma^{0}}$0.002 $0$0.007 -0.001 -0.009 1.000 0.032 0.033
$\alpha_{\Lambda}$ 0.018 0.011 -0.011 -0.051 -0.083 0.032 1.000 0.980
$\bar{\alpha}_{\Lambda}$0.013 0.027 0.004 -0.067 -0.083 0.033 0.980 1.000
Table 2: Values and uncertainties of the fit parameters and the $CP$ asymmetries $A^{\Sigma}_{CP}$ and $A^{\Lambda}_{CP}$, along with the previous measurements. The first and second uncertainties in this work are statistical and systematic, respectively, and those for the previous results are the total uncertainties. In the previous results, only the values reported in Ref. Ablikim et al. (2017) are cited for $\alpha_{J/\psi}$ and $\alpha_{\psi(3686)}$, as it gives the most precise measurement for $\alpha_{J/\psi}$ and the only measurement for $\alpha_{\psi(3686)}$.
Parameter This work Previous results
$\alpha_{J/\psi}$ $-0.4133\pm 0.0035\pm 0.0077$ $-0.449\pm 0.022$ Ablikim et al. (2017)
$\Delta\Phi_{J/\psi}$(rad) $-0.0828\pm 0.0068\pm 0.0033$
$\alpha_{\psi(3686)}$ $0.814\pm 0.028\pm 0.028$ $0.71\pm 0.12$ Ablikim et al. (2017)
$\Delta\Phi_{\psi(3686)}$(rad) $0.512\pm 0.085\pm 0.034$
$\alpha_{\Sigma^{0}}$ $-0.0017\pm 0.0021\pm 0.0018$
$\bar{\alpha}_{\Sigma^{0}}$ ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}0.0021\pm 0.0020\pm 0.0022$
$\alpha_{\Lambda}$ ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}0.730\pm 0.051\pm 0.011$ ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}0.748\pm 0.007$ Workman et al. (2022)
$\bar{\alpha}_{\Lambda}$ $-0.776\pm 0.054\pm 0.010$ $-0.757\pm 0.004$ Workman et al. (2022)
$A^{\Sigma}_{CP}$ ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}(0.4\pm 2.9\pm 1.3)\times 10^{-3}$
$A^{\Lambda}_{CP}$ $(-3.0\pm 6.9\pm 1.5)\times 10^{-2}$ $(-2.5\pm 4.8)\times 10^{-3}$ Ablikim et al. (2022b)
The transverse polarizations (parallel or anti-parallel to $\hat{y}_{\Sigma^{0}}$) of the $\Sigma^{0}$ hyperon in the $J/\psi$ and $\psi(3686)$ decays are observed for the first time and are found to have opposite directions. The corresponding parameters are determined to be $\Delta\Phi_{J/\psi}=-0.0828\pm 0.0068\pm 0.0033$ rad and $\Delta\Phi_{\psi(3686)}=0.512\pm 0.085\pm 0.035$ rad for the $J/\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ and $\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ decays, respectively, which differ from zero with a significance of $12.2\sigma$ for the $J/\psi$ dataset and $7.4\sigma$ for the $\psi(3686)$ dataset. The $\Sigma^{0}$ polarizations can be illustrated through the moment $\mu$, as defined:
$\mu^{k}(\cos\theta_{\Sigma})$ $=$
$\frac{1}{N_{\rm total}}\sum^{N^{k}}_{i}(\sin\theta^{i}_{\Lambda}\sin\varphi^{i}_{\Lambda}\cos\theta^{i}_{p}$ $+\sin\theta^{i}_{\bar{\Lambda}}\sin\varphi^{i}_{\bar{\Lambda}}\cos\theta^{i}_{\bar{p}}),$ (4)
where $N_{\rm total}$ is the total number of events in the dataset, $N^{k}$ is the number of events in the $k$-th $\cos\theta_{\Sigma}$ bin, and $i$ is the $i$-th event in that bin. The expected angular dependence of the moment for the acceptance-corrected data is $(\alpha_{\Lambda}-\bar{\alpha}_{\Lambda})(1+\alpha_{\psi}\cos^{2}\theta_{\Sigma})P_{y}(\cos\theta_{\Sigma})/(18+6\alpha_{\psi})$. Comparing the data to the PHSP MC sample, as shown in Fig. 2, the polarizations of the $\Sigma^{0}$ are observed clearly.
Figure 2: Distributions of the moments $\mu$ versus $\cos\theta_{\Sigma}$ for $J/\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ and $\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$. The points with error bars are data, the red solid lines are the signal MC samples with input parameters fixed to the fit results, and the blue dashed lines represent the distributions without polarization from the PHSP MC samples.
[…]
$\begin{split}\rm{T1}^{k}&=\frac{1}{N_{\rm total}}\sum_{i}^{N^{k}}\left(\cos^{2}\theta_{\Sigma}n^{i}_{1,z}n^{i}_{2,z}-\sin^{2}\theta_{\Sigma}n^{i}_{1,x}n^{i}_{2,x}\right),\\ \rm{T2}^{k}&=\frac{1}{N_{\rm total}}\sum_{i}^{N^{k}}\cos\theta_{\Sigma}\sin\theta_{\Sigma}\left(n^{i}_{1,z}n^{i}_{2,x}-n^{i}_{1,x}n^{i}_{2,z}\right),\\ \rm{T3}^{k}&=\frac{1}{N_{\rm total}}\sum_{i}^{N^{k}}\cos\theta_{\Sigma}\sin\theta_{\Sigma}\left(n^{i}_{1,y}+n^{i}_{2,y}\right),\\ \rm{T4}^{k}&=\frac{1}{N_{\rm total}}\sum_{i}^{N^{k}}\left(n^{i}_{1,z}n^{i}_{2,z}-\sin^{2}\theta_{\Sigma}n^{i}_{1,y}n^{i}_{2,y}\right),\end{split}$ (5)
where $n_{1,x}=\cos\theta_{p}\sin\theta_{\Lambda}\cos\varphi_{\Lambda}$, $n_{1,y}=\cos\theta_{p}\sin\theta_{\Lambda}\sin\varphi_{\Lambda}$, $n_{1,z}=\cos\theta_{p}\cos\theta_{\Lambda}$, $n_{2,x}=\cos\theta_{\bar{p}}\sin\theta_{\bar{\Lambda}}\cos\varphi_{\bar{\Lambda}}$, $n_{2,y}=\cos\theta_{\bar{p}}\sin\theta_{\bar{\Lambda}}\sin\varphi_{\bar{\Lambda}}$, $n_{2,z}=\cos\theta_{\bar{p}}\cos\theta_{\bar{\Lambda}}$. Figure 3 shows the distributions of these moments versus $\cos\theta_{\Sigma}$ for the $J/\psi$ dataset, the ones for the $\psi(3686)$ dataset are shown in Ref. 4. The fit results are consistent with those distirbutions in data.
[…]
Table 3: Absolute systematic uncertainties (×$\times 10^{-3}$) for the measured parameters.
Source ($10^{-3}$) $\alpha_{J/\psi}$ $\Delta\Phi_{J/\psi}$ (rad) $\alpha_{\psi(3686)}$ $\Delta\Phi_{\psi(3686)}$ (rad) $\alpha_{\Sigma^{0}}$ $\bar{\alpha}_{\Sigma^{0}}$ $\alpha_{\Lambda}$ $\bar{\alpha}_{\Lambda}$ $A^{\Sigma}_{CP}$ $A^{\Lambda}_{CP}$
Tracking 0.0 0.2 10 13 0.9 1.5 1 1 0.6 0
$\gamma$ reconstruction 0.5 0.5 1 0 0.2 0.1 2 3 0.0 3
Λ/$\Lambda/\bar{\Lambda}$ reconstruction 6.2 0.2 24 30 0.7 0.4 6 3 0.2 8
4C kinematic fit 1.5 0.1 5 6 1.0 1.5 2 3 0.6 3
$\Sigma^{0}/\bar{\Sigma}^{0}$ mass window 0.2 3.2 2 7 0.8 0.2 7 6 0.8 8
Background estimation 4.3 0.4 10 6 0.6 0.1 6 6 0.6 9
Total 7.7 3.3 28 34 1.8 2.2 11 10 1.3 15
The sources of the systematic uncertainties are summarized in Table 3. The total systematic uncertainties of various parameters are obtained by summing the individual contributions in quadrature. The details are described as follows.
The control samples $J/\psi\to p\bar{p}\pi^{+}\pi^{-}$, $e^{+}e^{-}\to\gamma\mu^{+}\mu^{-}$, and $J/\psi\to pK^{-}\bar{\Lambda}+c.c.$ are used to estimate the uncertainties from the $p,~{}\pi$ tracking, $\gamma$ reconstruction, and Λ/$\Lambda/\bar{\Lambda}$ reconstruction Ablikim et al. (2023d), respectively. The efficiency differences between data and MC simulation for the control samples are used to re-weight the PHSP MC sample. The differences between the fit results with corrections and the nominal fit are taken as the systematic uncertainties.
For the 4C kinematic fit, some discrepancies are observed in the $\chi^{2}_{\rm 4C}$ distributions between data and the signal MC sample. Similar to Refs. Ablikim et al. (2023e, 2024), a data-driven method is used to estimate these effects on the final results. The PHSP MC samples are weighted according to the $\chi^{2}_{\rm 4C}$ distributions in data, and the differences between the fit results after weighting and the nominal results are taken as the systematic uncertainties of the 4C kinematic fit.
The mass window of $\Sigma^{0}/\bar{\Sigma}^{0}$ is ±$\pm 3\sigma$ wide around the known $\Sigma^{0}$ mass Workman et al. (2022), where $\sigma=4.5$ MeV/$c^{2}$ is the mass resolution of the reconstructed $\Sigma^{0}/\bar{\Sigma}^{0}$. We change the mass window to ±$\pm 2\sigma$ or ±$\pm 4\sigma$ to study the systematic uncertainties from the $\Sigma^{0}/\bar{\Sigma}^{0}$ mass requirement. The largest deviations from the nominal values are taken as the systematic uncertainties.
The systematic uncertainties from the background estimation are studied by including the background in the fit or not. In the nominal solution, the events in the data sideband region are used to subtract the background events. Here, an alternative solution is performed by regarding all the selected events in the data sample as signal events. The differences between the alternative and nominal fit results are taken as the systematic uncertainties.
In summary, this Letter presents the first measurements of the parity-violating decay parameters for the decays $\Sigma^{0}\to\Lambda\gamma$ and $\bar{\Sigma}^{0}\to\bar{\Lambda}\gamma$, $\alpha_{\Sigma^{0}}=-0.0017\pm 0.0021\pm 0.0018$ and $\bar{\alpha}_{\Sigma^{0}}=0.0021\pm 0.0020\pm 0.0022$, respectively. The extracted values of decay parameters allow for the strong-$CP$ test, $A^{\Sigma}_{CP}=(0.4\pm 2.9\pm 1.3)\times 10^{-3}$. These results are consistent with $P$ and $CP$ conservation, and will provide crucial constraints for new physics models. The decay parameters $\alpha_{\Lambda}$ and $\bar{\alpha}_{\Lambda}$, which are listed in Table 2, are measured independently, and the weak-$CP$ test is performed in the sub-decays of $\Lambda$ and $\bar{\Lambda}$. These results are in good agreement with the values obtained from the $J/\psi\to\Lambda\bar{\Lambda}$ and $J/\psi\to\Xi^{-(0)}\bar{\Xi}^{+(0)}$ analyses of BESIII Ablikim et al. (2022a, b, 2023a).
Furthermore, the spin polarizations of the $\Sigma^{0}$ hyperons with opposite directions in the $J/\psi$ and $\psi(3686)$ decays are observed for the first time. This phenomenon is also observed in the case of the $\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ decays Ablikim et al. (2020a), but not in the $\Xi^{-(0)}$ Ablikim et al. (2023a, b, 2022a, 2022c) hyperon pairs. Following Ref. Wu et al. (2021) with the measured parameters $\alpha_{\psi}$ and $\Delta\Phi_{\psi}$, the ratio of the D-wave and S-wave coupling constants $g_{D}/g_{S}$, the relative phase $\delta$ between the S-wave and D-wave, the percentage of the S-wave contribution $\Gamma_{S}/\Gamma_{\rm total}$, and the effective radius $r_{\rm eff}$ of the $\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ decay (where $\bar{L}=r_{\rm eff}\times p$, $\bar{L}$ and $p$ are the average orbital angular momentum and the relative momentum of the $\Sigma^{0}$ and $\bar{\Sigma}^{0}$ in $\psi$ C.M. system) are obtained for the first time, as shown in Table 4 together with the parameters for $\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ calculated in Ref. Wu et al. (2021). What is particularly intriguing is that, as speculated in Ref. Wu et al. (2021), the $\delta$ difference between $J/\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ and $\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ is about the same as the value between $J/\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ and $\psi(3686)\to\Sigma^{+}\bar{\Sigma}^{-}$ decays Wu et al. (2021), and both are equal to approximately $\pi$. If this is not a coincidence, the underlying mechanism should be investigated in the future. These results are crucial to understand the decay dynamics of the charmonium states and the production mechanism of the $\Sigma^{0}-\bar{\Sigma}^{0}$ pairs and provide useful information for hunting down excited nucleon resonances Wu et al. (2021).
Table 4: The parameters of the $\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ (this work) and $\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ (from Ref. Wu et al. (2021)) processes, including the ratio between the coupling constants of D-wave and S-wave ($g_{D}/g_{S}$), the relative phase between S-wave and D-wave ($\delta$), the ratio between partial width of S-wave and total width ($\Gamma_{S}/\Gamma_{\rm total}$), and the effective radius ($r_{\rm eff}$). The first and second uncertainties in this work are statistical and systematic, respectively, and the ones for $\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ are the total uncertainties.
Mode $g_{D}/g_{S}$ (GeV-2) $\delta$ (rad) $\Gamma_{S}/\Gamma_{\rm total}$ (%) $r_{\rm eff}$ (fm)
$J/\psi\to\Sigma^{0}\bar{\Sigma}^{0}$ $0.1217\pm 0.0015\pm 0.0028$ $2.947\pm 0.017\pm 0.012$ $95.23\pm 0.11\pm 0.21$ $0.0191\pm 0.0005\pm 0.0008$
$\psi(3686)\to\Sigma^{0}\bar{\Sigma}^{0}$ $0.123\pm 0.008\pm 0.006$ $-0.28\pm 0.07\pm 0.04$ $82.7\pm 1.9\pm 1.4$ $0.049\pm 0.005\pm 0.004$
$J/\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ Wu et al. (2021) $0.171\pm 0.006$ $2.67\pm 0.04$ $90.9\pm 0.6$ $0.0362\pm 0.0024$
$\psi(3686)\to\Sigma^{+}\bar{\Sigma}^{-}$ Wu et al. (2021) $0.097\pm 0.009$ $-0.33\pm 0.10$ $88.3\pm 2.0$ $0.033\pm 0.006$
Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. The authors would like to extend thanks to Prof. Jusak Tandean, Dr. Shu-Ming Wu and Dr. Jia-Jun Wu for useful discussion and helpful advice. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2020YFA0406300, 2020YFA0406400; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11735014, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12165022, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12342502, 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; Yunnan Fundamental Research Project under Contract No. 202301AT070162; German Research Foundation DFG under Contracts Nos. 455635585, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; The Knut and Alice Wallenberg Foundation, Sweden; The Swedish Foundation for International Cooperation in Research and Higher Education (STINT); U. S. Department of Energy under Contract No. DE-FG02-05ER41374.